# Signature Kernel Scoring Rule as Spatio-Temporal Diagnostic for Probabilistic Forecasting

### v1

@@ Abstract
Modern weather forecasting has increasingly transitioned from numerical weather prediction (NWP) to data-driven machine learning forecasting techniques. While these new models produce probabilistic forecasts to quantify uncertainty, their training and evaluation may remain hindered by conventional scoring rules, primarily MSE, which ignore the highly correlated data structures present in weather and atmospheric systems. This work introduces the signature kernel scoring rule, grounded in rough path theory, which reframes weather variables as continuous paths to encode temporal and spatial dependencies through iterated integrals. Validated as strictly proper through the use of path augmentations to guarantee uniqueness, the signature kernel provides a theoretically robust metric for forecast verification and model training. Empirical evaluations through weather scorecards on WeatherBench 2 models demonstrate the signature kernel scoring rule’s high discriminative power and unique capacity to capture path-dependent interactions. Following previous demonstration of successful adversarial-free probabilistic training, we train sliding window generative neural networks using a predictive-sequential scoring rule on ERA5 reanalysis weather data. Using a lightweight model, we demonstrate that signature kernel based training outperforms climatology for forecast paths of up to fifteen timesteps.
@@ 1 Introduction
Weather variables exhibit strong spatial and temporal dependencies and are subject to non-stationary behaviour, with large-scale shifting due to seasonal patterns and climate trends. This variability makes uncertainty quantification essential, which is addressed by probabilistic forecasting. Using an ensemble method, the forecasting model makes a set of predictions, each with slightly perturbed weather model or initial conditions, producing a discrete approximation to the full forecast probability density function. Since 2022, new data-driven machine learning weather prediction approaches (MLWP) boosted by GPU acceleration have seen tremendous development. A variety of data-driven modelling approaches now demonstrate higher resolution, quicker computation, and greater performance on conventionally used forecasting scores than their physics-based NWP counterparts (Pathak et al., 2022; Keisler, 2022; Lam et al., 2023; Chen et al., 2023; Bi et al., 2023; Nguyen et al., 2023; Bouallègue et al., 2024; Lang et al., 2024a).
@@
However, the area of weather forecasting is noticeably limited in the justification provided for the scoring rules in use. In current literature, very few papers are using scores suggested by meteorological agencies, particularly the European Centre for Medium-Range Weather Forecast (ECMWF), despite many using their IFS and AIFS models as references (ECMWF, 2022). None have considered if specific weather variables, such as rainfall or temperature, require uniquely designed scoring rules, as opposed to general scoring rules or scores for other phenomena. A further effect is that there is no standard in scoring rule usage, with each developed model reporting different performance metrics. As a result, there is a marked need to accurately compare between the multitude of new data driven forecasting approaches. Moreover, most existing papers use certain scoring rules because of their simplicity, particularly versions of mean absolute error (MAE) and mean squared error (MSE), which have become the convention as more papers are published (Xu et al., 2024). Significantly, these scores are for deterministic point forecasts instead of ensemble forecasts. Many forecasting models are trained deterministically and then manipulated into an ensemble. This choice reflect a trade-off between computation efficiency and performance, particularly as the traditional method for generative probabilistic training is adversarial, which is highly computationally expensive (Scher and Messori, 2021).
In this work, we would propose and study the efficacy of a new diagnostic based on the signature kernel score as a standardising benchmark for probabilistic forecasting tasks involving spatio-temporal dependency, relevant to the particular problem of weather forecasting. Recently, Pacchiardi et al. (2024) proposed a framework of predictive-sequential (prequential) scoring rule to train probabilistic forecasting models, which was adapted for training AIFS in recent works (Lang et al., 2024b). Using our proposed signature kernel score and an extension of the prequential framework via sliding window generative models, we would show that we can train generative models to provide excellent multi-step probabilistic forecasting in future global weather prediction scenarios. Our contributions are:
@@
  • We demonstrate evaluation with inclusion on two weather scorecards.
  • We test our scoring rule on ERA5 reanalysis weather data, demonstrating its potential for training probabilistic weather forecasting models.
Section 2 introduces the necessary background of scoring rules and discusses their use in modern weather forecasting literature. The signature kernel scoring rule is established and shown to meet the theoretical requirements of a strictly proper scoring rule. Section 3 demonstrates the diagnostic utility of the signature kernel score with weather scorecards, considering the existing weather models of the IFS ens, Neural GCM, and FUXI. Section 4 contains the simulation results of training on ERA5 reanalysis data, considering performance via standard metrics and visual behaviour across ensemble members and path lengths.
Specific details of implementation are found in the appendix. Explanation of evaluation metrics used are in Appendix A. Objective functions and implementation approaches for evaluation and training are found in Appendix B and Section 4. Comments on numerical stability and Weatherbench data are found in Appendix C and D.
@@ 2 A new diagnostic for spatio-temporal probabilistic forecasting
In this section, we would introduce a new diagnostic which is a strictly proper probabilistic scoring rule based on signatures of path that accounts for both temporal and spatial dependencies to adequately address the critical needs of probabilistic weather forecasting. First we would provide a brief review of scoring rules commonly used as diagnostic for probabilistic forecasting, in Section 2.2 we introduce the signature of a path, discuss how to extend them to weather data observed on discrete time points in Section 2.3 and some common augmentations to signatures of a path suitable for the purpose in Section 2.4. Finally, in Section 2.5 we would propose the signature kernel scoring rule.
2.1 Commonly used Scoring Rules for diagnostic purpose
@@
A strictly proper scoring rule further ensures the equality holds if and only if $P=Q$, i.e. $Q$ is the unique minimiser. The inequality direction can be changed depending on whether the goal is maximisation or minimisation of the score, with minimisation taken to be the objective throughout the rest of this work without further reference. Strict propriety is necessary to prevent attempts at gaming the scoring rule by hedging, i.e. providing predictions that are not truthful, but are produced to achieve better scores. Serious issues potentially arise when a forecaster isn’t encouraged to report their true belief, and in the context of truthfully predicting the weather, improper scoring rules are a detriment to model performance. During model training, strictly proper scoring rules naturally lend themselves as loss functions by guaranteeing that the true distribution is the unique expected minimizer of the score. However, this theoretical result only holds in expectation, and in practice we can only consider the empirical score, which is the average score across observations for the predicted distribution $P$:
$\hat{S}_{N}(P)=\frac{1}{N}\sum^{N}_{i=1}S(P,y_{i})$
As a result, there is no guarantee of a unique minimiser in the loss landscape. Nonetheless, by the law of large numbers, the empirical score converges to the expected score as the number of i.i.d. samples $y_{i}\sim P$ goes to infinity:
$\hat{S}_{N}(P)\xrightarrow{\text{a.s.}}\mathbb{E}_{Y\sim Q}\left[S(P,Y)\right]$
Therefore, strictly proper scoring rules should result in models minimising predictive uncertainty, and resulting in forecasts approaching the true observation distribution. We will revisit the empirical score again when we discuss the extensions required for implementation and the prequential scoring rule.
Table 1 displays a summary analysis of the scoring rules in use among eight of the leading ML data-driven weather forecasting models between 2022-2024, determined by inclusion in the Google WeatherBench 2 framework (Rasp et al., 2020). These notably include the models from the researchers at Google DeepMind (Graphcast, GCM), Nvidia (FourCastNet FCN), and Huawei (Pangu-Weather PGW).
Scoring Rule / Paper FCN GNN GraphCast GCM FuXi PGW PGW Op. Stormer AIFS Probabilistic a b RMSE MAE ACC CRPS SSR Activity Bias RQE Calibration ROC MPE SEEPS a Diagnostics not evaluated on the ensemble version of Stormer. b AIFS is in ongoing development for ensemble forecasting.
Table 1: Summary of included scoring rules across eight significant ML-based weather prediction models (as identified by WeatherBench 2). PGW Op. refers to the operational version of Pangu-Weather. Nonstandard scores include ACC (Anomaly Correlation Coefficient), RQE (Relative Quantile Error), MPE (Mean Position Error), and SEEPS (Stable Equitable Error in Probability Space).
@@
Root Mean Squared Error (RMSE) is the most commonly used diagnostic, used on geopotential at 500 hPa and temperature at 850 hPa in every paper reviewed. These variables hold significant meteorological justification for underlying weather systems, and are the primary variables of focus for this paper (ECMWF, 2024a). There are limitations associated with training on RMSE, notably that it results in overly smooth forecast distributions by penalising bold predictions relevant for quantifying extreme weather risk (Bouallègue et al., 2024). The forecast activity, should therefore be a highly significant metric to contextualise differences in RMSE, but is only used in two papers reviewed. As mentioned in WeatherBench, RMSE is chosen primarily because of its ease to compute, and not due to a greater significance in the context of weather prediction (Rasp et al., 2020). Along with undesirable penalising behaviour, there is no spatial or time dependency involved. RMSE measures the average magnitude of differences between predictions and observations, without capturing any correlation or structure within the data. This issue means RMSE is unadapted to the specific problem of climate prediction, despite its ease and widespread use. Moreover, despite a majority of models being probabilistic, most evaluation focuses on the accuracy of the ensemble mean point forecast, especially in relation to RMSE. Consequently, forecast uncertainty is under-represented in the analysis.
@@ 2.2 Signature of a Continuous Path
A d-dimensional path $X$ is defined as a continuous mapping from an interval $[a,b]$ to $\mathbb{R}^{d}$, written as $X:[a,b]\rightarrow\mathbb{R}^{d}$ (Chevyrev and Kormilitzin, 2016). We often consider paths parametrised by time, with $X_{t}=\{X_{t}^{1},X_{t}^{2},...,X_{t}^{d}\}$ representing the position in space through time with $t\in[a,b]$. Each $X^{i}$ is likewise a path, with $X^{i}:[a,b]\rightarrow\mathbb{R}^{1}$. A common form of time dependent path is simply defined as $X_{t}=\{X_{t}^{1},X_{t}^{2}\}=\{t,f(t)\}$, which is a generalised time series consisting of successive measures in one dimension over an interval. In this section, we consider smooth paths which are infinitely continuously differentiable, i.e. $C^{\infty}$ (Chevyrev and Kormilitzin, 2016).
@@
Early iterations are easy to interpret, with the first order integrals being the range of the path in a specific dimension, resulting in $d$ first order terms. We can extend to second order iterations, where we consider any two dimensions of the path $i_{1},i_{2}\in\{1,...,d\}$:
$S(X)^{i_{1}.i_{2}}_{a,b}=\int_{a<s<b}S(X)^{i_{1}}_{a,s}dX^{i_{2}}_{s}=\int_{a<r<s<b}dX^{i_{1}}_{r}X^{i_{2}}_{s}$
$S(X)^{i,i}_{a,b}=\frac{(X^{i}_{b}-X^{i}_{a})^{2}}{2}=\frac{(S(X)^{i}_{a,b})^{2}}{2}$ (1)
However, non-diagonal terms, referred to as cross terms, with $i_{1}\neq i_{2}$ for second order terms are more complicated, relating to the Lévy area of the two-dimensional path. The Levy area is the signed area between a path and a chord created between its endpoints, represented in Figure 1 (Chevyrev and Kormilitzin, 2016) with areas $A_{-}$ and $A_{+}$, denoted by:
@@
Figure 1: Representation of the signed Lévy area of a two-dimensional path. $A_{+}$ and $A_{-}$ denote the positive and negative respective areas between the path (red) and chord (blue). $S^{1,2}$ and $S^{2,1}$ denote the second order terms equal to the area between the path and their perpendicular enclosure (black).
We finally extend iteration to the k-iterated integral, for any collection of indexes $i_{1},...,i_{k}\in\{1,...,d\}$, we define:
These higher order terms no longer have simple geometric interpretations. However, we can extend the results found in Equations 1 and 2 using the shuffle product identity (Ree, 1958). This identity shows that we can relate the product of any two signature terms $S(X)_{a,b}^{i_{1},\ldots,i_{k}}$ and $S(X)_{a,b}^{j_{1},\ldots,j_{m}}$ as the sum of $S(X)^{K}_{a,b}$ terms, which depend only on indexes ($i_{1},\ldots,i_{k}$) and ($j_{1},\ldots,j_{m})$. This collection of new indexes, $K$, is obtained by permuting both index sets together while keeping their respective ordering, referred to as shuffling and denoted by . Given a permutation $\sigma$ of the set $i_{1},\ldots,i_{k},j_{1},\ldots,j_{m}$}, $\sigma$ is a $(k,m)$-shuffle if $\sigma^{-1}(i_{1})<\ldots<\sigma^{-1}(i_{k})$ and $\sigma^{-1}(j_{1})<\ldots<\sigma^{-1}(j_{m})$. We can therefore define the shuffle product identity:
$S(X)^{i_{1},...,i_{k}}_{a,b}\cdot S(X)^{j_{1},...,i_{m}}_{a,b}=\sum_{K\in\{i_{1},...,i_{k}\}\shuffle\{j_{1},...,i_{m}\}}S(X)^{K}_{a,b}$
Finally, the signature of path X, denoted by $S(X)_{a,b}$ on the set $\left[a,b\right]$ is the infinite collection of all iterated integrals of X, which can be written as an infinite sequence of real numbers:
[…]
The zeroth order term is considered to be 1, by convention. This infinite signature captures all of the sequential dependencies in the path, with each order of term capturing increasingly complex relationships across space over time. The properties of the signature can be compared to moments in statistics. While moments capture characteristics of a distribution, the signature characterises the temporal and structural data within a path. As the moments determine a distribution, albeit under some conditions, the signature likewise uniquely characterises a path, up to translation and time parametrization, which will be addressed in an upcoming subsection. Given the properties of capturing path behaviour, with terms of the signature being good candidates for characteristic features, the full signature contains all the information we want to extract from a path for a machine-learning algorithm to understand the behaviour.
2.3 Extension to Discrete Weather Data
The previous section considered signatures defined for smooth paths. However, in our context, we would like to view our discrete time forecasts and observations as paths. Additionally, with handling real weather data, it is necessary to consider how irregularly sampled or partially observed data affects the result. This requires extending our definitions of the signature from paths to data streams consisting of discrete points. We consider a $d$-dimensional data stream, $X\in\mathbb{R}^{d}$ for $N$ points with an ordered set of times, $t_{0}<t_{1}<\dots<t_{N}$, as follows:
$X=\{\left(X^{1}_{t_{i}},\dots,X^{d}_{t_{N}}\right)\}^{N}_{i=0}$
$\hat{X}_{t}=X_{t_{i}}+\frac{t-t_{i}}{t_{i+1}-t_{i}}(X_{t_{i+1}}-X_{t_{i}})$
Each linear interpolation is piecewise infinitely continuously differentiable $C^{\infty}$ and satisfies the requirements of a smooth path. The overall path is the result of concatenating each interpolated section, resulting in a continuous $C^{0}$ path for any $t\in\left[t_{0},t_{N}\right]$ (Chevyrev and Kormilitzin, 2016). We now revisit the calculation of the signature by considering signatures of each interpolated linear path on their respective interval $t_{i}$, $t_{i+1}$]. Letting $x$ be the straight line defined on $t_{i}$, $t_{i+1}$], we define the signature:
$S(X)_{t_{i},t_{i+1}}^{j_{1},\ldots,j_{k}}=\frac{1}{k!}\prod^{k}_{l=1}\left(x_{j_{l}}(t_{i+1})-x_{j_{l}}(t_{i})\right)$
$S(X)_{t_{a},t_{c}}^{j_{1},\ldots,j_{k}}=\sum_{l=0}^{k}S(X)_{t_{a},t_{b}}^{j_{1},j_{2},\ldots,j_{l-1}}\cdot S(X)_{t_{b},t_{c}}^{j_{l},j_{l+1},\ldots,j_{k}}$
$S(X)^{1}_{t_{a},t_{c}}=S(X)^{1}_{t_{a},t_{b}}+S(X)^{1}_{t_{b},t_{c}}$
$S(X)^{4}_{t_{a},t_{c}}=S(X)^{4}_{t_{a},t_{b}}+S(X)^{3}_{t_{a},t_{b}}\otimes\ldots$
Now the signature is defined for the region $t_{a}$, $t_{c}$], and can be extended to further points $t_{d}$ with step by step concatenation following the same method. We now have a method to consider weather data and our associated forecasts as paths which can be appropriately transformed into a signature, capturing the dependencies of our chronological sequence across the latitude-longitude grid. However, there are further extensions which need to be taken to augment the path to improve the performance of the signature.

### v2

# Signature Kernel Scoring Rule: A Spatio-Temporal Diagnostic for Probabilistic Weather Forecasting

@@ Abstract
Modern weather forecasting has increasingly transitioned from numerical weather prediction (NWP) to data-driven machine learning forecasting techniques. While these new models produce probabilistic forecasts to quantify uncertainty, their training and evaluation may remain hindered by conventional scoring rules, primarily MSE, which are designed for single time point predictions and ignore the highly correlated data structures present in weather behaviour. This work introduces the signature kernel scoring rule to the domain of weather forecasting, which reframes weather variables as continuous paths to encode temporal and spatial dependencies through iterated integrals. Validated as strictly proper through the use of path augmentations to guarantee uniqueness, the signature kernel provides a theoretically robust metric for forecast verification and model training. Empirical evaluations through weather scorecards on WeatherBench 2 models demonstrate the signature kernel scoring rule’s high discriminative power and unique capacity to capture path-dependent interactions. Following previous demonstration of successful adversarial-free probabilistic training, we train sliding window generative neural networks using a predictive-sequential scoring rule on ERA5 reanalysis weather data. Using a lightweight model, we demonstrate that signature kernel based training outperforms climatology for forecast paths of up to fifteen timesteps.
@@ 1 Introduction
Weather variables exhibit strong spatial and temporal dependencies and are subject to non-stationary behaviour, with large-scale shifting due to seasonal patterns and climate trends. This variability makes uncertainty quantification essential, which is addressed by probabilistic forecasting. Using an ensemble method, the forecasting model makes a set of predictions, each with slightly perturbed weather model or initial conditions, producing a discrete approximation to the full forecast probability density function. Since 2022, new data-driven machine learning weather prediction approaches (MLWP) boosted by GPU acceleration have seen tremendous development. A variety of data-driven modelling approaches now demonstrate higher resolution, quicker computation, and greater performance on conventionally used forecasting scores than their physics-based NWP counterparts (Pathak et al., 2022; Keisler, 2022; Lam et al., 2023; Chen et al., 2023; Bi et al., 2023; Nguyen et al., 2023; Bouallègue et al., 2024b; Lang et al., 2024a).
@@
However, the area of weather forecasting is noticeably limited in the justification provided for the scoring rules in use. In current literature, very few papers are using scores suggested by meteorological agencies, particularly the European Centre for Medium-Range Weather Forecast (ECMWF), despite many using their Integrated Forecasting System (IFS) and Artificial Intelligence Forecasting System (AIFS) models as references (ECMWF, 2022b). A further effect is that there is no standard in scoring rule usage, with each developed model reporting different performance metrics. As a result, there is a marked need to accurately compare the multitude of new data driven forecasting approaches. Moreover, most existing papers use certain scoring rules because of their simplicity, particularly versions of mean absolute error (MAE) and mean squared error (MSE), which have become the convention as more papers are published (Xu et al., 2024). Furthermore, these scores only evaluate marginal forecast distributions at individual time steps, offering no assessment of whether a model correctly captures the temporal and spatial dependency structure over long forecast horizons. Significantly, many forecasting models are trained deterministically and then manipulated into an ensemble. This choice reflects a trade-off between computation efficiency and performance, particularly as the traditional method for generative probabilistic training is adversarial, which is highly computationally expensive (Scher and Messori, 2021).
In this work, we propose and study the efficacy of a new diagnostic based on the signature kernel score, which natively operates over entire forecast paths, as a standardising benchmark for probabilistic forecasting tasks involving spatio-temporal dependency, relevant to the particular problem of weather forecasting. The signature kernel was first developed in 2019 as a similarity measure for paths, grounded in rough path theory, and later optimised for efficient computation in 2021 (Kiraly and Oberhauser, 2019) (Salvi et al., 2021). We build on the theoretical framework of the signature kernel as a score, established in 2023, to consider long time step weather forecasting evaluations (Issa et al., 2023). The application of the signature kernel score to the high-dimensional domain of global weather prediction, where complex correlation structures stand to be taken advantage of by path curvature characteristics, remains a significant and unexplored challenge. Recently, (Pacchiardi et al., 2024) proposed a framework of predictive-sequential (prequential) scoring rule to train probabilistic forecasting models, which was adapted for training AIFS in recent works (Lang et al., 2024b). Using the signature kernel score and an extension of the prequential framework via sliding window generative models, we show that we can train generative models to provide excellent multi-step probabilistic forecasting in future global weather prediction scenarios. Our contributions are:
@@
  • We illustrate diagnostic performance with evaluation on two weather scorecards.
  • We demonstrate probabilistic training of our scoring rule on ECMWF reanalysis weather data.
Section 2 introduces the necessary background of scoring rules and discusses their use in modern weather forecasting literature. The signature kernel scoring rule is established and shown to meet the theoretical requirements of a strictly proper scoring rule. Section 3 demonstrates the diagnostic utility of the signature kernel score as a multiple time step evaluator with weather scorecards, considering the existing models of the IFS ENS (ensemble), Neural GCM (general circulation model), and FuXi Weather. Section 4 contains the simulation results of training on ERA5 reanalysis data, considering performance via standard metrics and visual behaviour across ensemble members and path lengths, along with a discussion on numerical stability.
All figures, scorecard generation code, results, trained models, and training regimes can be found in the following GitHub repository here. Explanation of evaluation metrics used are in Appendix A. The extended theory behind the signature kernel can be found in Appendix B, C, D and E. Extended objective functions and implementation approaches for evaluation and training used in Section 4 are found in Appendix F. An additional illustrative experiment on signature behaviour is shown in Appendix G. Comments on WeatherBench data are found in Appendix H. Figures of model architecture and additional training results are found in Appendix I and J. Further discussion of extensions and limitations of the work is found in Appendix K.
@@ 2 A new diagnostic for spatio-temporal probabilistic forecasting
In this section, we introduce a new diagnostic, which is a strictly proper probabilistic scoring rule based on the signature of a path that accounts for both temporal and spatial dependencies to adequately address the needs of probabilistic weather forecasting. First, we provide a brief review of scoring rules commonly used as diagnostic for probabilistic forecasting. In Section 2.2, we introduce the signature of a path. We discuss the extension to weather data observed on discrete time points in Section 2.3 and some common augmentations to signatures of a path suitable for the purpose in Section 2.4. Finally, in Section 2.5, we propose the signature kernel scoring rule.
2.1 Commonly used Scoring Rules for Diagnostic Purposes
@@
A strictly proper scoring rule further ensures the equality holds if and only if $P=Q$, i.e. $Q$ is the unique minimiser. The inequality direction can be changed depending on whether the goal is maximisation or minimisation of the score, with minimisation taken to be the objective throughout the rest of this work without further reference. Strict propriety is necessary to prevent attempts at gaming the scoring rule by hedging, i.e. providing predictions that are not truthful, but are produced to achieve better scores. Serious issues potentially arise when a forecaster isn’t encouraged to report their true belief, and in the context of truthfully predicting the weather, improper scoring rules are a detriment to model performance. During model training, strictly proper scoring rules naturally lend themselves as loss functions by guaranteeing that the true distribution is the unique expected minimiser of the score.
@@
Table A.1 (Appendix) displays a summary analysis of the scoring rules in use among eight of the leading ML data-driven weather forecasting models between 2022-2024, determined by inclusion in the Google WeatherBench 2 framework (Rasp et al., 2020). Root Mean Squared Error (RMSE) is the most commonly used diagnostic, used on geopotential at 500 hPa and temperature at 850 hPa in every paper reviewed. These variables hold significant meteorological justification for underlying weather systems, and are the primary variables of focus for this paper (ECMWF, 2024a). There are limitations associated with training on RMSE, notably that it results in overly smooth forecast distributions by penalising bold predictions relevant for quantifying extreme weather risk (Bouallègue et al., 2024b). The forecast activity should therefore be a highly significant metric to contextualise differences in RMSE, but is only used in two papers reviewed. As mentioned in WeatherBench, RMSE is chosen primarily because of its ease to compute, and not due to a greater significance in the context of weather prediction (Rasp et al., 2020). Along with undesirable penalising behaviour, there is no spatial or time dependency involved. RMSE measures the average magnitude of differences between predictions and observations, without capturing any correlation or structure within the data. Moreover, despite a majority of models being probabilistic, most evaluation focuses on the accuracy of the ensemble mean point forecast, especially in relation to RMSE. Consequently, forecast uncertainty is under-represented in the analysis.
@@ 2.2 Signature of a Continuous Path
A d-dimensional path $X$ is defined as a continuous mapping from an interval $[a,b]$ to $\mathbb{R}^{d}$, written as $X:[a,b]\rightarrow\mathbb{R}^{d}$ (Chevyrev and Kormilitzin, 2025). We often consider paths parametrised by time, with $X_{t}=\{X_{t}^{1},X_{t}^{2},...,X_{t}^{d}\}$ representing the position in space through time with $t\in[a,b]$. Each $X^{i}$ is likewise a path, with $X^{i}:[a,b]\rightarrow\mathbb{R}^{1}$. A common form of time dependent path is simply defined as $X_{t}=\{X_{t}^{1},X_{t}^{2}\}=\{t,f(t)\}$, which is a generalised time series consisting of successive measures in one dimension over an interval. In this section, we consider smooth paths which are infinitely continuously differentiable, i.e. $C^{\infty}$ (Chevyrev and Kormilitzin, 2025).
@@
Early iterations are easy to interpret, with the first order integrals being the range of the path in a specific dimension, resulting in $d$ first order terms. We can extend to second order iterations, relating to the encompassed area of the path, where we consider any two dimensions of the path $i_{1},i_{2}\in\{1,...,d\}$:
$S(X)^{i_{1},i_{2}}_{a,b}=\int_{a<s<b}S(X)^{i_{1}}_{a,s}dX^{i_{2}}_{s}=\int_{a<r<s<b}dX^{i_{1}}_{r}X^{i_{2}}_{s}$
We extend iteration to the k-iterated integral, for any collection of indexes $i_{1},...,i_{k}\in\{1,...,d\}$, we define:
These higher order terms no longer have simple geometric interpretations. See appendix B for a further geometric intuition of lower order terms. The signature of path X, denoted by $S(X)_{a,b}$ on the set $\left[a,b\right]$ is the infinite collection of all iterated integrals of X, which can be written as an infinite sequence of real numbers:
@@
The zeroth order term is considered to be 1, by convention. This infinite signature captures all of the sequential dependencies in the path, with each order of term capturing increasingly complex relationships across space over time. The properties of the signature can be compared to moments in statistics. While moments capture characteristics of a distribution, the signature characterises the temporal and structural data within a path. As the moments determine a distribution, albeit under some conditions, the signature likewise uniquely characterises a path, up to translation and time parametrisation. Given the properties of capturing path behaviour, with terms of the signature being good candidates for characteristic features, the full signature contains all the information we want to extract from a path for a machine-learning algorithm to understand the behaviour. See Appendix C for the extension to discrete data.